What is the change in internal energy of a system if the system (i) absorbs 58 J of heat and does 58 J of work? (ii) absorbs 125 J of heat and does 687 J of work? (iii) evolves 280 cal of heat and has 1.25 kJ of work done on it?
Answer #1

(i) absorbs 58 J of heat. So, Q = +58 J

does 58 J of work. So, W = -58 J

use:

delta U = Q + W

= 58 J - 58 J

= 0

Answer: 0

(ii) absorbs 125 J of heat. So, Q = +125 J

does 687 J of work. So, W = -687 J

use:

delta U = Q + W

= 125 J - 687 J

= -562 J

Answer: -562 J

(iii) evolves 280 cal of heat.

So, Q = -280 cal

= -280*4.184 J

= -1171.52 J

= -1.17 KJ

has 1.25 kJ of work done on it.So, W = +1.25 KJ

use:

delta U = Q + W

= -1.17 KJ +1.25 KJ

= -0.08 KJ

Answer: -0.08 KJ